# LLMs Cannot Reliably Judge (Yet?): A Comprehensive Assessment on the Robustness of LLM-as-a-Judge

### v1

@@ 3.2 Defender
We categorize defenses for LLM-as-a-Judge into two main types: prevention and detection. Prevention-based defenses aim to redesign the evaluation prompt or pre-process inputs to block adversarial manipulations from influencing the model’s judgments. Detection-based defenses focus on identifying whether the evaluation data has been tampered with or compromised.
@@
Defender’s Capability. The defender cannot control the attacker’s actions or modify the model post-release. In closed-source settings, the defender may additionally implement dynamic defense mechanisms, including real-time detection of compromised inputs, instruction reinforcement to mitigate prompt injection attempts, and adaptive intervention during evaluation. Across both settings, the defender assumes that attackers can freely submit malicious candidate responses but cannot modify the judge’s internal model weights or core prompt instructions.
@@
Defender’s Goal. The defender’s goal is to safeguard the evaluation integrity of LLM-as-a-Judge systems by ensuring that judgments are based solely on the intended evaluation criteria, without being influenced by adversarial manipulations. Specifically, the defender aims to (i) prevent the execution of injected instructions through prompt design and input sanitization, and (ii) detect and mitigate compromised evaluation data when prevention fails.
An effective defense should preserve evaluation fidelity under both benign and adversarial conditions while minimizing utility degradation and false detections.
4 RobustJudge
4.1 Overview of RobustJudge
Figure 2: Overview of the RobustJudge workflow. RobustJudge is a framework to evaluate the robustness of LLM-as-a-Judge systems. It supports customizable configuration of key modules, including tasks, evaluation protocols (pointwise or pairwise), adversarial attacks (via Attacker Factory), defense techniques (via Defense Guards), judge prompt templates, and judge models.
RobustJudge aims to evaluate the robustness of LLM-as-a-Judge systems and generates quantitative metrics based on their judgment performance. As illustrated in Figure 2, RobustJudge follows a modular workflow that begins with selecting a benign input query from a specific tasks (e.g., translation, summarization, code generation)(§4.2). The response to this input, generated by a target LLM, is then processed through two parallel paths:
  • Benign path: The response is directly formatted using a specified Judge Prompt Template4.6) and sent to the LLM judges4.7) for evaluation.
  • Adversarial path: The same response is first manipulated by the Attacker Factory4.4), generating an adversarial variant. This response can optionally pass through the Defense Guard4.5) for mitigation, before being evaluated by the same LLM judge.
The LLM-as-a-Judge then evaluates both benign and adversarial responses, assigning scores (for pointwise evaluation protocol) or pairwise preferences (for pairwise evaluation protocol)(§4.3). In parallel, the responses are analyzed by a Content Quality Evaluator4.8), which provides content quality scores using automatic metrics, such as BLEURT for text tasks or CodeBLEU for code-related tasks. Finally, the results from both paths are consolidated in the Metrics module (§4.9), where RobustJudge quantifies robustness based on the judgments on the benign and adversarial responses and their content quality scores.
4.2 Tasks
In this paper, we evaluate LLM-as-a-Judge on 8 tasks. These tasks fall into 3 categories, e.g., text-focused, code-oriented and knowledge-intensive evaluation tasks. The details of these tasks are summarized in Table I.
@@ 4.2 Tasks
Text-focused Tasks. This category aims to evaluate the robustness of judgment on fundamental text processing tasks. Specifically, we include two tasks, e.g., machine translation and text summarization.
Machine Translation (T1). This task focuses on translating text between languages while maintaining meaning, fluency, and accuracy. We adopt 6 language pairs from FLORES-200 benchmark [41], including high-resource pairs such as Chinese-English, German-English, as well as the low-resource Yoruba-English, all evaluated in both translation directions.
Text Summarization (T2). Text Summarization distills lengthy documents into concise, factual summaries. We draw on English news articles from the CNN/DailyMail dataset [42]. Each sample in this dataset contains an original article and its reference summary.
@@
Code-oriented Tasks. Processing and generating code are essential capabilities for assessing the performance of both LLMs and LLM-as-a-Judge systems. To comprehensively evaluate judgment in the code domain, we include three representative tasks: code translation, code summarization, and code generation.
Code Translation (T3). Code Translation involves converting source code from one programming language to another while preserving its original functionality. For this evaluation, we use the Code-TransOcean dataset [42], focusing on the subset with C#-Java language pairs.
Code Summarization (T4). This task involves generating concise natural language descriptions that capture the functionality of given code snippets. For our evaluation, we utilize the Python code summarization subset from the CodeXGLUE benchmark  [43].
Code Generation (T5). The code generation task focuses on producing executable code based on a natural language specification or a function signature. Our evaluation is based on the samples from the CodeSearchNet dataset [44], specifically targeting Python function generation.
TABLE I: Judge tasks used in our dataset.
ID Category Task Name #Nums. Source Dataset
T1 Text Text Translation 30 Flores-200[41]
T2 Text Summarization 20 CNN/DailyMail[42]
T3 Code Code Translation 30 CodeTransOcean[45]
T4 Code Generation 20 CodeSearchNet[44]
T5 Code Summarization 20 CodeXGLUE[43]
T6 Knowledge Mathematics 20 LiveBench[46]
T7 Logical Reasoning 20 LiveBench[46]
T8 Knowledge Recall 28 MMLU-Pro[47]
@@
Knowledge-intensive Tasks. We extend the evaluation of the LLM-as-a-Judge system’s robustness in knowledge-intensive scenarios. Following the setup in JudgeBench [8], we select three core areas, i.e., logical reasoning, mathematical problem solving, and professional knowledge recall.
logical Reasoning (T6). We incorporate benchmarks from Big-Bench Hard [48] and classic puzzles (e.g., the Zebra Puzzle) to examine the judgment robustness on abstract, compositional, and logical inference skills. These tasks measure the model’s ability to perform multi-step reasoning and handle complex constraint-based problems.
Mathematical Problem Solving (T7). Our mathematical evaluation suite comprises problems from AMC12 and USAMO competitions, including topics such as algebra, geometry, and combinatorics. These problems assess the model’s ability to perform numerical reasoning and formal proof strategies.
@@
Professional Knowledge Recall (T8). We employ MMLU-Pro [47], an enhancement of the original MMLU [49]. MMLU-Pro contains college-level, multiple-choice questions across 14 professional disciplines, e.g., Physics, Chemistry, and Law.
4.3 Evaluation Protocol
$s=M(P_{1},r)$ (1)
Pairwise Evaluation. This protocol compares two candidate responses and determines which one is preferred in the given context. The model outputs a preference for either response $r_{a}$ or $r_{b}$, depending on their relative quality. Formally, this can be expressed as:
$p=M(P_{2},r_{a},r_{b})$ (2)
where $P_{2}$ is the pairwise prompt template, and $r_{a}$ and $r_{b}$ are the two responses under evaluation, $p\in\{r_{a},r_{b}\}$. Prior research on LLM-as-a-Judge systems [5] has shown that the input order of the candidate responses can affect the evaluation outcome. To account for this bias, evaluations are typically conducted using both input orders, with the results defined as:
$p_{+}=M(P_{2},r_{a},r_{b}),\quad p_{-}=M(P_{2},r_{b},r_{a})$ (3)
4.4 Attacker Factory
@@
Attacker factory is responsible for generating adversarial responses $r_{m}$ by applying various attack techniques designed to probe the vulnerabilities of LLM-as-a-Judge systems. These attacks are broadly categorized into heuristic-based and optimization-based methods, depending on how adversarial attacks are constructed. A summary of heuristic-based attack methods is listed in Table VIII (appendix), and optimization-based attacks can be found in Table IX (appendix).
@@
Heuristic-based Attacks. These attacks are based on manually crafted prompts or linguistic manipulations that exploit LLM weaknesses in instruction following, context interpretation, or reasoning flow. Unlike optimization-based approaches, these attacks aim to alter the judgment outcome through direct prompt manipulations. Formally, a heuristic-based attack is defined as:
$r_{m}=A(r)$ (4)
@@
where $r$ denotes the original benign response, and $A$ is the transformation applied by the attack. Our evaluation includes the following 8 heuristic-based attacks: Naive Attack (H1) [30], Escape Characters (H2) [31], Context Ignoring (H3) [32, 39], Fake Completion (H4) [40], Fake Reasoning (H5) [12], Combined Attack (H6) [14], Empty (H7) and Long-Suffix (H8).
Optimization-Based Attacks. These attack methods employ automated search or optimization algorithms to progressively manipulate response $r$. They iteratively modify candidate responses based on judge scores or internal signals. This process is formalized as:
$r_{m}^{(i+1)}=A\left(r_{m}^{(i)}\right),\quad\text{for }i=0,1,\ldots,n$ (5)
where $r_{m}^{(0)}=r$ is the initial benign response, and $n$ is the number of iterations. The 7 optimization-based attacks in our evaluation include: AdvEval (O1) [10], PAIR (O2) [15], TAP (O3) [16], Cheating (O4) [11], GCG (O5) [33], AutoDAN (O6) [34], and Greedy (O7) [9].
4.5 Defense Guard
@@
To evaluate the impact of defense techniques against adversarial attacks for LLM-as-a-Judge systems, we incorporate a Defense Guard module in RobustJudge. Inspired by PromptBench [14], this module includes both prevention-based and detection-based strategies, which are lightweight and applicable without modifying the LLM judges. Further technical details are provided in Appendix  IX.
Prevention-based Defenses. These methods aim to enhance the actual input $I$ to the judge models by modifying the prompt $P$ and/or the response $r$ to prevent attacks from taking effect. This can be expressed as:
$I=D(P,r)$ (6)
where $D$ denotes the transformation function, such as re-tokenizing inputs or reformatting the prompt. We evaluate the following prevention methods: Retokenization (D1) [17], Delimiters (D2) [40], Sandwich Prevention (D3) [35], Instructional Prevention (D4) [50].
Detection-based Defenses. These methods aim to identify and filter out adversarial inputs before they reach the judge model. If an input is flagged as adversarial based on a predefined metric $f(\cdot)$, the evaluation is skipped. Formally:
$I=(P,r),\quad\text{iff }f(P,r)>\tau$ (7)
where $f(\cdot)$ outputs a scalar confidence score and $\tau$ is a detection threshold. We consider three detection strategies: PPL (D5) [37], Windowed PPL (D6) [17], Naive LLM-based (D7) [18].
4.6 Judge Prompt Template
@@
In the context of LLM-as-a-Judge, a variety of judge prompt templates have been developed to guide LLMs in evaluating model outputs. To explore how these templates influence the robustness of LLM-as-a-Judge systems, we begin by collecting 3 widely adopted judge prompt templates, e.g., Vanilla Prompt[51], Arena-Hard Prompt[52], and Google Vertex Prompt[53] from JudgeBench [8]. We evaluate their performance under various adversarial attacks to examine their contribution to the robustness of LLM-as-a-Judge systems.
@@
We hypothesize that the choice of judge prompt template plays a critical role in shaping the judge system’s robustness. Therefore, identifying more effective templates is essential for improving their robustness. To explore this further, we extend our analysis to include 6 additional commonly used prompt templates: PandaLM, Prometheus 2, JudgeLM, Auto-J, Skywork, and ChatEval. All these 9 instances collectively form a representative set of judge prompt templates.
@@
TABLE II: Summary of key components of judge prompt template.
@@
This process continues until convergence, yielding a prompt configuration that maximizes evaluation accuracy and robustness. We present the coordinate ascent algorithm in Algorithm 1.
@@
0:  Prompt components $\mathcal{C}=\{C_{1},C_{2},\dots,C_{n}\}$, initial configuration $\mathbf{c}^{(0)}$, evaluation function $\mathcal{E}(\cdot)$
0:  Optimized prompt configuration $\mathbf{c}^{*}$
1:  Initialize $\mathbf{c}^{(0)}$ (e.g., randomly or using a known baseline)
2:  Set $t\leftarrow 0$
3:  repeat
4:     for each component $C_{i}\in\mathcal{C}$ do
5:        Fix all components except $C_{i}$
6:        Search for the best $c_{i}^{(t+1)}$ that maximizes $\mathcal{E}(\mathbf{c})$
7:        Update: $\mathbf{c}^{(t+1)}\leftarrow\mathbf{c}^{(t)}$ with $C_{i}=c_{i}^{(t+1)}$
8:     end for
9:     $t\leftarrow t+1$
10:  until convergence or maximum iterations reached
11:  return $\mathbf{c}^{(t)}$ as $\mathbf{c}^{*}$
4.7 LLM Judges
We consider 12 LLM-as-a-Judge systems with diverse characteristics across multiple dimensions: backbone architecture, model size, fine-tuning status, reasoning capability, and open-source availability.
Specifically, the evaluated LLM judges include proprietary systems such as GPT-4o and PAI-Judge, as well as open-source models built on LLaMA, Mistral, and Qwen backbones. Among them, several models (e.g., JudgeLM, PandaLM, Auto-J, Prometheus 2) have been explicitly fine-tuned for judgment tasks, while others (e.g., LLama-3.1, Openchat-3.5, Qwen2.5) remain zero-shot evaluators. We also include Ds-R1, a reasoning-oriented judge system built on Ds-V3, to explore whether enhanced reasoning capabilities influence judgment robustness and behavior. The features of the LLM judges are provided in Table XI (appendix).
4.8 Content Quality Evaluator
Our evaluation metrics (detailed in §4.9) are designed to capture not only the judgment accuracy of the LLM-as-a-Judge systems but also the content quality drift of the evaluated responses. To quantify this drift, we introduce a content quality evaluator module $Q$, which assesses how closely a target response aligns with a reference response. Formally, the content quality score $S_{e}$ is defined as:

### v2

@@ III-C Defender
We categorize defenses into two types: detection-based and prevention-based. Detection-based defenses identify whether evaluation data has been compromised. Prevention-based defenses redesign evaluation prompts or preprocess inputs to block adversarial manipulations from influencing judgments.
@@
Defender Capabilities. In closed-source settings, defenders may implement dynamic defense mechanisms, including real-time detection of compromised inputs, instruction reinforcement to mitigate prompt injection, and adaptive intervention during evaluation.
@@
Defender Goals. The defender aims to safeguard evaluation integrity by ensuring that judgments are based solely on the intended criteria, without any adversarial influence. Specifically, the defender seeks to: (i) prevent execution of injected instructions through prompt design and input sanitization, and (ii) detect and mitigate compromised data when prevention fails. An effective defense should preserve evaluation fidelity under both benign and adversarial conditions while minimizing utility degradation and false detections.
Figure 2: The framework consists of five components: (§IV-A) Dataset Construction supporting diverse task types (text, vision, code, knowledge); (§IV-B) Attacker Factory implementing heuristic and optimization-based attacks; (§IV-C) Defense Guard deploying detection and prevention mechanisms; (§IV-D) LLM Judge with customizable prompts and evaluation metrics (pairwise/scoring); and (§V) Result Analysis assessing attack effectiveness, defense performance, prompt impact, and real-world vulnerabilities.
IV RobustJudge
In this section, we introduce the construction process of RobustJudge, as illustrated in Figure 2.
IV-A Dataset Construction
TABLE I: Overview of our benchmark dataset.
ID Category Task Name Source Dataset
T1 Text Text Translation FLORES-200 flores200
T2 Text Summarization CNN/DailyMail cnn
T3 Code Code Translation CodeTransOcean CodeTransOcean
T4 Code Summarization CodeXGLUE CodeXGLUE
T5 Code Generation CodeSearchNet codesearchnet
T6 Knowledge Logical Reasoning LiveBench livebench
T7 Mathematics LiveBench livebench
T8 Knowledge Recall MMLU-Pro mmlupro
We systematically construct a benchmark to assess the robustness of LLM-as-a-Judge systems across three task categories: text-focused, code-oriented, and knowledge-intensive tasks, with a detailed breakdown in Table I. For each task, we first generate a high-quality response to a given prompt [^2] We use OpenChat-3.5-0106 in this paper., which is then fed into the attacker factory to systematically generate adversarial variants.
IV-A1 Text-Focused Tasks
@@
This category evaluates the robustness of judges on standard natural language processing tasks, specifically machine translation and text summarization.
Machine Translation (T1). We select $30$ sentences from the FLORES-200 flores200 test set, spanning six language pairs, including high-resource languages (e.g., Chinese–English, German–English) and low-resource languages (e.g., Yoruba–English). Evaluations are performed bidirectionally for each pair to assess the judge’s robustness in evaluating translation quality, including fluency, accuracy, and semantic fidelity, under adversarial perturbations.
Text Summarization (T2). We use $20$ articles from the CNN/DailyMail cnn test split to evaluate the judge’s ability to assess the factual consistency and conciseness of summaries when subjected to adversarial manipulations.
IV-A2 Code-Oriented Tasks
@@
Code comprehension and generation are essential capabilities for evaluating LLM and LLM-as-a-Judge systems. We include three representative tasks: code translation, code summarization, and code generation.
Code Translation (T3). We select $30$ samples from the CodeTransOcean dataset CodeTransOcean , covering six widely used programming languages (C++, Python, Go, Java, C, and C#), to evaluate the judge’s robustness in assessing cross-language code translation under adversarial conditions.
Code Summarization (T4). We use $30$ samples from the Python subset of CodeXGLUE CodeXGLUE to assess whether adversarial inputs compromise the judge’s reliability in evaluating code-to-documentation generation.
Code Generation (T5). Based on $30$ samples from CodeSearchNet codesearchnet , we evaluate the judge’s robustness in assessing Python function generation from natural language specifications or function signatures under adversarial attacks.
IV-A3 Knowledge-Intensive Tasks
@@
We evaluate judge robustness in knowledge-intensive scenarios across three domains: logical reasoning, mathematical problem solving, and professional knowledge recall, following the JudgeBench framework judgebench .
Logical Reasoning (T6). We use $56$ samples from Big-Bench Hard bigbench and classic logic puzzles (e.g., the Zebra Puzzle) to test the judge’s robustness in evaluating abstract, compositional, and constraint-based reasoning under adversarial attacks.
Mathematical Problem Solving (T7). We select $98$ problems from AMC12 and USAMO competitions spanning algebra, geometry, and combinatorics to assess the judge’s robustness in evaluating mathematical solutions against adversarially crafted inputs.
@@
Professional Knowledge Recall (T8). We employ $134$ samples from MMLU-Pro mmlupro , an enhanced version of MMLU mmlu , containing college-level multiple-choice questions across 14 professional disciplines (e.g., Physics, Chemistry, and Law).
IV-B Attacker Factory
@@
Within RobustJudge, the attacker factory generates adversarial responses $r_{m}$ through diverse techniques designed to expose vulnerabilities in LLM-as-a-Judge systems. We categorize these attacks into heuristic-based and optimization-based methods.
IV-B1 Heuristic-Based Attacks
@@
Heuristic-based attacks exploit weaknesses in instruction following, context interpretation, and reasoning by injecting manually crafted prompts or linguistic perturbations. Formally, such attacks can be expressed as:
$r_{m}=f_{\text{LLM}}(r\oplus h),$ (1)
where $r$ is the original response, $h$ denotes a set of adversarial perturbations, $\oplus$ indicates insertion or concatenation, and $r_{m}$ is the resulting adversarial response.
@@
We consider eight heuristic-based attacks: Naive Attack (H1) naive_2 , Escape Characters (H2) injection , Context Ignoring (H3) context_ignore_1 ; context_ignore_2 , Fake Completion (H4) fake_completion , Fake Reasoning (H5) judgeDeceiver , Combined Attack (H6) prompt_bench , Empty Attack (H7), and Long-Suffix Attack (H8).
IV-B2 Optimization-Based Attacks
Optimization-based attacks construct adversarial suffixes by iteratively optimizing an objective function with respect to the judge model. The optimization process can be expressed as:
$s^{*}=\arg\min_{s}\;\mathcal{L}_{\text{judge}}\!\left(f_{\text{LLM}}(r\oplus s),y\right),$ (2)
where $\mathcal{L}_{\text{judge}}$ denotes the loss function defined by the judge model, $y$ is the target label, $s$ is the adversarial suffix, and $s^{*}$ is the optimized suffix.
These attacks differ in their design of the objective function and optimization strategy. Some methods directly minimize classification loss (e.g., AdvEval (O1) adv ), while others optimize for task-specific objectives such as consistency with adversarial prompts (PAIR (O2) pair , TAP (O3) tap ) or explicitly craft misleading reasoning paths (Cheating (O4) cheating ). Gradient-based approaches such as GCG (O5) gcg and AutoDAN (O6) autodan iteratively refine suffixes through differentiable optimization, whereas Greedy Attack(O7) op2 employs heuristic optimization.
IV-C Defense Guard
@@
To evaluate defense effectiveness against adversarial attacks on LLM-as-a-Judge systems, we incorporate a Defense Guard module in RobustJudge. Inspired by PromptBench prompt_bench , this module includes both detection-based and prevention-based strategies that can be applied without modifying the underlying judge models.
IV-C1 Detection-Based Defenses
A detector $G(\cdot)$ evaluates the prompt-response pair $(P,r)$ by computing a detection metric $M$ and comparing it against a threshold $\tau$:
$G(P,r)=\begin{cases}\text{Safe},&M\leq\tau,\\ \text{Unsafe},&M>\tau,\end{cases}$ (3)
where $M$ represents the detection score (e.g., perplexity-based or anomaly score), and $\tau$ is the decision threshold. Inputs flagged as unsafe are blocked before reaching the judge model.
We evaluate three detection strategies: Perplexity Filter  decetor , Windowed Perplexity  baseDefense , and Naive LLM-based Detector  naive_llm_defense .
IV-C2 Prevention-Based Defenses
Prevention-based defenses augment the judge prompt by appending defensive content $a$ to the original prompt $P$. Formally, the defended prompt is expressed as:
$P^{\prime}_{\text{judge}}=P_{\text{judge}}\oplus a,$ (4)
where $\oplus$ denotes concatenation and $a$ is a defensive augmentation designed to mitigate adversarial influence.
We evaluate four prevention methods: Re-tokenization  baseDefense , Delimiters  fake_completion , Sandwich Prevention  sandwich , and Instructional Prevention  instructiondefense .
IV-D LLM Judges
This section details the RobustJudge testing pipeline for LLM-as-a-Judge [^3] All judge models evaluated in this work are summarized in Table IV.
IV-D1 Judge Prompt Template
In the context of LLM-as-a-Judge, numerous prompt templates have been proposed to instruct models in evaluating the quality of their output. While such systems are expected to exhibit consistent safety performance across prompts when assessing the same task, we observe a pronounced prompt instability: their robustness against adversarial attacks varies significantly with the choice of prompt.
@@ IV-D1 Judge Prompt Template
To explore how these templates influence the robustness of LLM-as-a-Judge systems, we begin by collecting 3 widely adopted judge prompt templates, e.g., Vanilla PromptVanilla , Arena-Hard Prompthard , and Google Vertex Promptgoogle from JudgeBench judgebench . RobustJudge evaluate their performance under various adversarial attacks to examine their contribution to the robustness of LLM-as-a-Judge systems.
@@
We hypothesize that the choice of judge prompt template plays a crucial role in shaping the robustness of the judge system. Therefore, identifying more effective templates is essential for improving their robustness. To explore this further, we extend our analysis to include 6 additional commonly used prompt templates: PandaLM, Prometheus 2, JudgeLM, Auto-J, Skywork, and ChatEval. These 9 instances collectively form a representative set of judge prompt templates.
@@
TABLE II: Summary of key components in judge prompt templates. RS: Role Specification; EI: Evaluation Instruction; IAG: In-context Answer Guideline; EC: Evaluation Criteria; ER: Evaluation Requirement; RF: Response Format. Symbols: ✓= included, ✗= not included.
@@
This process continues until convergence, yielding a prompt configuration that maximizes evaluation accuracy and robustness. We present the coordinate ascent algorithm Algorithm 1 to facilitate subsequent comparison.
@@
0: Components $\mathcal{C}=\{C_{1},\ldots,C_{n}\}$, init config $\mathbf{c}^{(0)}$, eval func $\mathcal{E}(\cdot)$, max iter $T_{\max}$
0: Optimized config $\mathbf{c}^{*}$
1:Init: $\mathbf{c}^{(0)}$, $t\leftarrow 0$
2:repeat
3:  for $i=1$ to $n$ do
4:   Fix $\mathbf{c}_{-i}^{(t)}$ {All except $c_{i}$}
5:   $c_{i}^{(t+1)}\leftarrow\arg\max_{c_{i}}\mathcal{E}(\mathbf{c}_{-i}^{(t)},c_{i})$
6:   Update $\mathbf{c}^{(t+1)}$
7:  end for
8:  $t\leftarrow t+1$
9:until $\|\mathbf{c}^{(t)}-\mathbf{c}^{(t-1)}\|<\epsilon$ or $t\geq T_{\max}$
10:return $\mathbf{c}^{*}=\mathbf{c}^{(t)}$
IV-D2 Content Quality Evaluator
Our evaluation framework (detailed in §IV-D3) is designed to capture two distinct aspects of adversarial attacks on LLM-as-a-Judge systems: (1) direct manipulation of the judgment mechanism, and (2) indirect effects through content quality changes in evaluated responses. To disentangle these effects, we introduce a content quality evaluator module $Q$ that measures content-level drift between clean and attacked responses.
The evaluator serves as a differential measurement tool rather than an absolute quality assessor. Its purpose is to quantify how much an attack alters the semantic or functional characteristics of responses, enabling us to distinguish whether judgment shifts arise from compromised judgment logic or from genuine content changes that should naturally influence scores. The content quality score $S_{e}$ is formally defined as: